15. Fu Jewelers purchased 4,000,000 of the outstanding 12,000,000 shares of Harris & Associates. At the time of the acquisition, the book value of Harris's net assets equal their fair market value. Harris declared dividends of $300,000 during the year. How will Fu record this transaction?

a. Fu will increase the investment account by $100,000. b. Fu will increase Dividend Revenue by $100,000. c. Fu will increase Dividend Revenue by $300,000. d. Fu will decrease the investment account by $100,000.16. The following summary data for Amirkhani Construction Co.s (ACC) relates to Job 1227, which was completed in 2019, are p
Answer #1

Solution 15:

% of holding of FU Jewellers in Harris & Associates = 4000000 / 12000000 = 33.33%

Therefore FU is having significant influence in Harris & Associates.

Share of dividend received by FU = $300,000 * 33.33333% = $100,000

Fu will decrease the investment account by $100,000 for receipt of dividend.

Hence option d is correct.

Solution 16:

Computation of % of completion, revenue recognition and Gross Profit
Year Actual cost incurred (A) Total cost incurred till date (B) Total estimated Cost ( C ) % of completion (D) (B/C) Contract Price
(E)
Total Revenue to be recoganized (F) (E * D) Revenue for current period (G) Gross Profit (H)
2018 $180,000 $180,000 $375,000 48.00% $450,000 $216,000 $216,000 $36,000
2019 $200,000 $380,000 $380,000 100.00% $450,000 $450,000 $234,000 $34,000

Hence option c is correct.